A 0.2025 Sample consisting of only Bacl2 and Kcl required 20.25ml of 0.1200M Agno3 solution for the quantitative precipitation of chloride. Calculate the %Ba and %K in the sample.


1
Expert's answer
2021-03-15T09:09:34-0400

2AgNO3 + BaCl2 = 2AgCl + Ba(NO3)2

CM = n/V

n = CM x V

n (AgNO3) = 0.1200 x 0.02025 = 0.00243 mol

n (BaCl2) = 1/2 x n (AgNO3) = 0.00243/2 = 0.001215 mol

n = m/M

M (BaCl2) = 208.23 g/mol

m (BaCl2) = n x M = 0.001215 x 208.23 = 0.25 g


The mass of BaCl2 in the sample is 0.25 g. Mass of the sample is required to calculate the %Ba and %K in the sample.